translate the statement into an algebraic expression or equation.
The sales for this month S2 are 20% greater than the sales from last month S1.

Answer by ikleyn(52835)

.

    S2 = S1 + 0.2*S1;   

or

    S2 = (1 + 0.2)*S1,

or

   S2 = 1.2*S1.